A manufacturer of electronics products is considering entering the telephone equipment business. It estimates that if it were to begin making wireless telephones, its short-run cost function would be as follows:

Q(thousands)        AVC          AC                MC


         9                  41.10       52.21           30.70

        10                 40.00       50.00           30.10

       11               39.10       48.19            30.10

       12                 38.40       46.73            30.70

      13                 37.90      45.59             31.90

      14                 37.60        44.74             33.70

     15                37.50         44.17             36.10

      16                37.60         43.85             39.10

     17               37.90        43.78             42.70

     18                38.40       43.96             46.90

     19                 39.10       44.36              51.70

     20               40.00       45.00               57.10

a. Plot the average cost, average variable cost, marginal cost, and price on a graph

b. explain the reasons for the shape or curve of AC,AVC and MC.

c. Suppose the average wholesale price of a wireless phone is currently $50. Do you think this company should enter the market? Explain.

d. what is the optimum level of production for the producer? Indicate the amount of profit (or loss) earned by the firm at the optimum level of production.


1
Expert's answer
2021-10-27T19:03:10-0400

a.


b.

p=mc therefore the firm should enter the market.

c.

"p=50\\mc=51.7\\Q=19\\TR=19\times 50=950\\AC=44.36\\AC=\frac{total\space cost}{output}\\44.36=\frac{TC}{19}\\TC=842.84\\profit=TR-TC\\=950-842.84=107.26"